OpenStudy (anonymous):

Translate into an inequality. Janice and Leonard together would like to save more than $3,500. they have save $ 425 saved so far. Janice can save $120 more than twice what Leonard can save each month. how much money will they each need to save to reach their goal.

OpenStudy (anonymous):

do you need this inequality solved as well?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so janice ( J ) and leonard ( L ) want to save more than 3500 but they want to do it together. so we can represent this by: J + L > 3500

OpenStudy (anonymous):

J=120+2L from the third sentence

OpenStudy (anonymous):

right now J+L=450

OpenStudy (anonymous):

brainfart

OpenStudy (anonymous):

Hint: System of equations, napervillian.

OpenStudy (anonymous):

L = 450 - J

OpenStudy (anonymous):

^what he said

OpenStudy (anonymous):

You may want to elaborate on it, as it might not be very obvious as to what the next step is. :P

OpenStudy (anonymous):

ok